OpenStudy (anonymous):

1.8q-3.6-4.8q=-4.0q-2.9

OpenStudy (ddcamp):

Start by combining like terms: 1.8q - 3.6 - 4.8q = 1.8q - 4.8q - 3.6 = -3.0q - 3.6 Then isolate the variable (q) -3.0q - 3.6 = -4.0q - 2.9 -3.0q + 4.0q - 3.6 = -2.9 1.0q - 3.6 = -2.9 q = .7

OpenStudy (anonymous):

how did u get the .7?

OpenStudy (ddcamp):

-2.9 + 3.6

OpenStudy (ddcamp):

So I added 3.6 to both sides

OpenStudy (anonymous):

oooh ok i see now lol thank u

OpenStudy (anonymous):

is z/9=-7 -63